OpenStudy (anonymous):

What is d/dx of cos^2 (x^3) ?

OpenStudy (anonymous):

\[d/dx [\cos^2(x^3)]\]

OpenStudy (anonymous):

Use chain rule and try :)

OpenStudy (anonymous):

Ohh wooaah [x3] You know you love me, I know you care Just shout whenever, and I'll be there You are my love, you are my heart And we will never ever ever be apart Are we an item? Girl, quit playing We're just friends, what are you saying? Say there's another and look right in my eyes My first love broke my heart for the first time And I was like... Baby, baby, baby oooh Like baby, baby, baby nooo Like baby, baby, baby oooh I thought you'd always be mine (mine) Baby, baby, baby oooh Like baby, baby, baby nooo Like baby, baby, baby oooh I thought you'd always be mine (mine) Oh, for you I would have done whatever And I just can't believe we ain't together And I wanna play it cool, but I'm losin' you I'll buy you anything, I'll buy you any ring And I'm in pieces, baby fix me And just shake me 'til you wake me from this bad dream I'm going down, down, down, down And I just can't believe my first love won't be around And I'm like Baby, baby, baby oooh Like baby, baby, baby nooo Like baby, baby, baby oooh I thought you'd always be mine (mine) Baby, baby, baby oooh Like baby, baby, baby nooo Like baby, baby, baby oooh I thought you'd always be mine (mine) [Ludacris:] Luda! When I was 13, I had my first love, There was nobody that compared to my baby and nobody came between us or could ever come above She had me going crazy, oh, I was star-struck, she woke me up daily, don’t need no Starbucks. She made my heart pound, it skipped a beat when I see her in the street and at school on the playground but I really wanna see her on the weekend. She knows she got me dazing cause she was so amazing and now my heart is breaking but I just keep on saying... Baby, baby, baby oooh Like baby, baby, baby nooo Like baby, baby, baby oooh I thought you'd always be mine (mine) Baby, baby, baby oooh Like baby, baby, baby nooo Like baby, baby, baby oooh I thought you'd always be mine (mine) I'm gone (Yeah Yeah Yeah, Yeah Yeah Yeah) Now I'm all gone (Yeah Yeah Yeah, Yeah Yeah Yeah) Now I'm all gone (Yeah Yeah Yeah, Yeah Yeah Yeah) Now I'm all gone (gone, gone, gone...) I'm gone

OpenStudy (anonymous):

Please explain how 2 differentiate using chain rule....i can never do it correcily

OpenStudy (anonymous):

\[d/dx= 2cosx^{3}(-sinx^{3})3x^{2}\]

OpenStudy (anonymous):

can u explain the chain rule?

OpenStudy (anonymous):

i mean thats not correct

OpenStudy (anonymous):

-3 x^2 sin(2 x^3)

OpenStudy (anonymous):

sorry the 9 is supposed 2 be a parenthesis. it should be -6x^2 (sinx^3) *cos(x^3)

OpenStudy (anonymous):

\[d/dx(\cos^{2}(x^{3}))=2\cos(x^{3})(d/dx(\cos(x^{3})\]

OpenStudy (anonymous):

\[2\cos(x^{3})(-\sin(x^{3})(d/dx(x^{3}))\]

OpenStudy (anonymous):

\[-2\cos(x^{3})\sin(x^{3})3x^{2}\]

OpenStudy (anonymous):

\[-6x^{2}\sin(x^{3})\cos(x^{3})\]

OpenStudy (anonymous):

why did the2cos(s^3) turn negative to -2cos(x^3)

OpenStudy (anonymous):

because the (-) on the sine can be factored out

OpenStudy (anonymous):

please expalin the gist of the chain rule. taht is what u used right?

OpenStudy (anonymous):

Chain rule should be applied like. Eg Differentiate y =cos(sin(tan(cot(x)))) Now we should assume sin(tan(cot(x)) as t y = cos t dy/dt = -sin t * (dt/dx) Now continue like that

OpenStudy (anonymous):

so for my problem it would be like y = cos x and dy/dx = -sin (x^3)....idk chain rule is confusing...can u esplain it using the problem i posted

OpenStudy (anonymous):

@help_with_math , see this video. It should help you :) http://www.khanacademy.org/math/calculus/v/the-chain-rule

OpenStudy (anonymous):

ok thanks